Hi. Any help would be greatly appreciated.
Find the magnitude and direction of the resultant of the following 3 forces in 2 dimensions.
Force 1 10N at 30 degrees to the positive x axis
Force 2 20N at 120 degrees to the positive x axis
Force 3 30N at 240 degrees to the positive x axis.
Thank you

Answer by Alan3354(69443)

Find the magnitude and direction of the resultant of the following 3 forces in 2 dimensions.
Force 1 10N at 30 degrees to the positive x axis
Force 2 20N at 120 degrees to the positive x axis
Force 3 30N at 240 degrees to the positive x axis.
==========
Do you mean all 3 forces at the same time?
If the 3 forces act at the same point:
Force 1 10N at 30 degrees to the positive x axis
x component = 10*cos(30) = 5sqrt(3)
y component = 10*sin(30) = 5
---
Force 2 20N at 120 degrees to the positive x axis
x component = 20*cos(120) = -10
y component = 20*sin(120) = 10sqrt(3)
---
Force 3 30N at 240 degrees to the positive x axis.
x component = 30*cos(240) = -15
y component = 30*sin(240) = -15sqrt(3)
=================================
Sum of x components = -25 + 5sqrt(3) =~ -16.34
Sum of y components = 5 - 5sqrt(3) =~ -3.66
========
Angle = atan(y/x) in Q3 =~ 192.63 degs
R = sqrt(x^2 + y^2) =~ 16.74N